# Regularization of the ensemble Kalman filter using a non-parametric, non-stationary spatial model

### v3

@@ Appendix E The most localized kernel in the stationary case
where it is straightforward to see that $\operatorname{\mathbb{E}}\widetilde{\xi}_{\ell m}\widetilde{\xi}_{\ell^{\prime}m^{\prime}}^{*}=\sigma_{\ell}^{2}\,\delta_{\ell}^{\ell^{\prime}}\delta_{m}^{m^{\prime}}$. The (real-valued) spectrum of the process is $f_{\ell}=\sigma_{\ell}^{2}$.
The process $\xi_{\rm statio}(x)$ can be modeled as the convolution of the kernel
$u(\rho)=\sum\frac{2\ell+1}{4\pi}\sigma_{\ell}P_{\ell}(\cos\rho)$ (45)
@@
Appendix F Identifiability of the model
@@
where $Q$ is the half-bandwidth of the processes $\sigma_{\ell}(x)$ as functions of location $x$. Note that since $\sigma_{-\ell}(x)=\sigma_{\ell}(x)$ (see D), we have $\widetilde{\sigma}_{-\ell,q}=\widetilde{\sigma}_{\ell q}$. Since $\sigma_{\ell}(x)$ is real-valued, we have $\widetilde{\sigma}_{\ell,-q}=\widetilde{\sigma}_{\ell q}^{*}$.
@@
Let us substitute the expansion Eq. (49) into Eq. (42) and change the summation variable $q$ to $n=\ell+q$555 In this Appendix (in contrast to the rest of the paper), we denote $\ell_{\rm max}$ by $L$ to simplify the notation. :
@@
so that we can uniquely restore $\widetilde{\sigma}_{Lq}$ for all $q$. To realize this, take $q=-Q$ in Eq. (52) (the lower right corner of the summation area in Fig. 8). We obtain $\widetilde{B}_{N,L-Q}=\widetilde{\sigma}_{LQ}\,\widetilde{\sigma}_{L,-Q}^{*}$. Recalling that $\widetilde{\sigma}_{\ell,-q}=\widetilde{\sigma}_{\ell q}^{*}$, we have $\widetilde{B}_{N,L-Q}=\widetilde{\sigma}_{LQ}^{2}$, which implies that $\widetilde{\sigma}_{LQ}$ is uniquely determined by $\widetilde{B}_{N,L-Q}$, up to sign. By the tight bandwidth assumption, $\widetilde{\sigma}_{LQ}\neq 0$, therefore, we recover $\widetilde{\sigma}_{Lq}$ for all $|q|<Q$ from Eq. (52) (the right edge of the summation parallelogram in Fig. 8), again, up to sign. The ambiguity in the sign disappears when it comes to $\widetilde{\sigma}_{L0}$, which is positive because $\sigma_{\ell}(x)>0$ due to the positive-definiteness constraint (see item 2 in section 2.8 and D). Thus, $\widetilde{\sigma}_{Lq}$ are uniquely recovered for all $q$ from the model covariances.
@@
Then, we consider $\widetilde{\xi}_{N-1}$ and realize that it, again, has just one term, $\widetilde{\alpha}_{L-1}\,\widetilde{\sigma}_{L-1,L-1+Q}$, besides the term that contains $\widetilde{\sigma}_{L,L+Q-1}$, which has already been recovered. This allows us to repeat the above process and recover $\widetilde{\sigma}_{L-1,L-1+q}$ for all $q\in[-Q,Q]$. And so on, we uniquely recover all non-zero $\widetilde{\sigma}_{\ell q}$, and therefore all spectral functions $\sigma_{\ell}(x)$, from the set of spectral covariances $\widetilde{B}_{n\bar{n}}$. This completes the proof that the non-stationary spatial convolution model on the circle is unique provided that the convolution kernel $u(x,z)$ is real-valued, is an even function of $z$, and the half-bandwidth $Q$ of $\sigma_{\ell}(x)$ is such that $\widetilde{\sigma}_{\ell Q}\neq 0$ for any $\ell$. On ${\mathbb{S}}^{2}$, the same reasoning is valid (not shown).
Appendix G Local stationarity
The notion of local stationarity has been defined differently by different authors (most often for processes on the real line). The general idea is that a locally stationary process can be approximated by a stationary process locally, i.e., in the vicinity of any point in time (Mallat et al., 1998). Starting from Dahlhaus (1997), the common approach is to use the “infill” asymptotics, which considers the process $\xi(t)$ in rescaled time $t/T$ with $t=0,1,\dots,T$ and $T\to\infty$. With this approach, effectively, just one segment of the non-stationary process is studied in increasingly sharper detail.
On compact manifolds like the circle or the sphere, this rescaling cannot be used because of their compactness. Here we outline a different approach to local stationarity with our spatial model. For simplicity, we consider the circular case (extensions to the spherical case are straightforward).
Recall that the non-stationary random process in question is defined as $\xi(x)=\sum\sigma_{\ell}(x)\,\widetilde{\alpha}_{l}\,\mathrm{e}^{\mathsf{i}\ell x}$ (Eq. (42)). As its non-stationarity comes from the dependencies of $\sigma_{\ell}$ on the location $x$, let us assume that these are caused by $\sigma_{\ell}(x)$ being random processes themselves. More specifically, we postulate that $\sigma_{\ell}(x)$ are, jointly, a vector-valued mean-square differentiable stationary random process. Then, we obtain a spatial non-stationarity pattern by considering the conditional distribution of the process $\xi$ given all $\sigma_{\ell}(\cdot)$. We denote this conditioning by $\xi\,|\,\sigma$.
$\operatorname{\mathbb{E}}\left(\frac{\partial\sigma_{\ell}}{\partial x}\right)^{2}=\frac{{\mathsf{Var}}\,\sigma_{\ell}}{\Lambda_{\ell}^{2}}.$ (53)
If $\min_{\ell}(\Lambda_{\ell})=\Lambda>0$, we call $\Lambda$ the non-stationarity length scale.
Next, we define the characteristic length scale $L$ of the process $\xi(x)$ from
$L^{2}=\frac{\sum\operatorname{\mathbb{E}}\sigma_{\ell}^{2}}{\sum\ell^{2}\operatorname{\mathbb{E}}\sigma_{\ell}^{2}},$ (54)
i.e., as the inverse width of the mean local spectrum.
Finally, we define the non-stationarity strength per wavenumber,
$S_{\ell}=\frac{{\mathsf{Var}}\,\sigma_{\ell}}{\operatorname{\mathbb{E}}^{2}\sigma_{\ell}},$ (55)
which quantifies the magnitude of the random component in the local spectrum, ${\mathsf{Var}}\,\sigma_{\ell}$, relative to its mean value squared, $\operatorname{\mathbb{E}}^{2}\sigma_{\ell}$. We let the overall non-stationarity strength be $S=\max_{\ell}S_{\ell}\in[0,1)$.
Having defined $\Lambda$, $L$, an $S$, we find conditions under which $\xi(x)$ is locally stationary in the following sense. For any point $x_{0}$, there is an approximating stationary random process $\zeta(x;x_{0})$ such that $\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}[(\xi(x)-\zeta(x;x_{0}))^{2}\,|\,\sigma]\ll\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}[\xi^{2}(x)\,|\,\sigma]$ whenever $|x-x_{0}|\leq L$. Formally, we can consider the limit $\Lambda/L\to\infty$, that is, the large $\Lambda/L$ asymptotics, but we prefer to specify in which sense $\Lambda$ can be considered “large” with respect to $L$ in a practical situation.
We define the approximating process as $\zeta(x;x_{0})=\sum\sigma_{\ell}(x_{0})\,\widetilde{\alpha}_{l}\,\mathrm{e}^{\mathsf{i}\ell x}$, that is, with the argument of $\sigma_{\ell}$ fixed at $x_{0}$. Then, the mean approximation error variance is
$\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}[(\xi(x)-\zeta(x;x_{0}))^{2}\,|\,\sigma]=\sum\operatorname{\mathbb{E}}(\sigma_{\ell}(x)-\sigma_{\ell}(x_{0}))^{2}=\\ \sum\operatorname{\mathbb{E}}\left(\frac{\partial\sigma_{\ell}}{\partial x}\right)^{2}(x-x_{0})^{2}\leq\left(\frac{L}{\Lambda}\right)^{2}\sum{\mathsf{Var}}\,\sigma_{\ell}.$ (56)
Here, the derivatives are taken at some points between $x_{0}$ and $x$. The inequality is due to Eq. (53), $\Lambda\leq\Lambda_{\ell}$, and $|x-x_{0}|\leq L$. Noting that $\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}[\xi^{2}(x)\,|\,\sigma]=\sum\operatorname{\mathbb{E}}\sigma_{\ell}^{2}=\sum(\operatorname{\mathbb{E}}^{2}\sigma_{\ell}+{\mathsf{Var}}\,\sigma_{\ell})$, we write down the relative approximation error variance:
$\frac{\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}[(\xi(x)-\zeta(x;x_{0}))^{2}\,|\,\sigma]}{\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}[\xi^{2}(x)\,|\,\sigma]}\leq\frac{\sum{\mathsf{Var}}\,\sigma_{\ell}}{\sum\operatorname{\mathbb{E}}^{2}\sigma_{\ell}+\sum{\mathsf{Var}}\,\sigma_{\ell}}\left(\frac{L}{\Lambda}\right)^{2}\leq S\left(\frac{L}{\Lambda}\right)^{2}.$ (57)
This equation reveals that the locally stationary approximation is precise whenever (i) $S\ll 1$ and $\Lambda\geq L$ or (ii) $\Lambda\gg L$ irrespective of $S$ (recall that $10\leq S<$ 1). The case of low non-stationarity strength, $S\ll 1$, is rather trivial because then the spatial variability in the local spectra is negligible so the process is almost stationary. The non-trivial local stationarity occurs when the the non-stationarity strength $S$ is non-negligible and the non-stationarity length scale $\Lambda$ is large enough for the upper bound of the approximation error given by Eq. (57) to be acceptably small.
If $\Lambda\gg L$, then it is easy to see that $L_{\xi}\approx L$, where $L_{\xi}$ is defined through $\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}(\xi_{x}^{2}\,|\,\sigma)={\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}(\xi^{2}\,|\,\sigma)}/{L_{\xi}^{2}}$ and $L$ is defined in Eq. (54). This justifies labeling $L$ as the “characteristic length scale of the process $\xi(x)$”.
Appendix H Bandpass filtering of a locally stationary signal
Here we consider a random process governed by the Locally Stationary Convolution Model and assess the approximation error involved in Eq. (8). For simplicity of presentation, we examine the circular case.
@@ Appendix H Bandpass filtering of a locally stationary signal
$H(\ell)=\kappa\left(\frac{\ell-\ell^{\rm c}}{d}\right),$ (58)
where $\kappa(\cdot)$ is a real-valued bell-shaped function of continuous argument such that $\kappa(0)=\max|\kappa(\cdot)|=1$ and $\max|\kappa^{\prime}(\cdot)|=1$, $\ell^{\rm c}$ is the band’s central wavenumber, and $d$ is the half-bandwidth (cf. Eq. (19)). Let $L_{\cal H}=1/d$ denote the width of the filter’s impulse response function. Apply the filter ${\cal H}$ to the locally stationary process
@@
$\varphi(x)=\sum H(n)\,\widetilde{\xi}_{n}\,\mathrm{e}^{\mathsf{i}nx}=\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\widetilde{\alpha}_{\ell}\,\mathrm{e}^{\mathsf{i}\ell x}\sum_{q=-Q}^{Q}H(\ell+q)\,\widetilde{\sigma}_{\ell q}\,\mathrm{e}^{\mathsf{i}qx}.$ (60)
As in G), we postulated that the processes $\widetilde{\sigma}_{\ell}(x)$ are, jointly, a mean-square-differentiable stationary vector-valued random process. This implies, in particular, that $\operatorname{\mathbb{E}}\widetilde{\sigma}_{\ell 0}=\operatorname{\mathbb{E}}\sigma_{\ell}$ and $\operatorname{\mathbb{E}}\widetilde{\sigma}_{\ell q}=0$ for $q\neq 0$.
Now, suppose that the spatial filter in question is selected such that $L_{\cal H}\ll\Lambda$ (here $\Lambda$ is the non-stationarity length scale defined in G). Then, $H(\ell+q)$ in Eq. (60) is much flatter than $\widetilde{\sigma}_{\ell q}$ (both considered as functions of $q$). Indeed, the squared width of $H(\ell)$ is $L_{\cal H}^{-2}$. The squared spectral width of $\widetilde{\sigma}_{\ell q}$ (as a function of $q$) can be assessed as the width of the function $\operatorname{\mathbb{E}}|\widetilde{\sigma}_{\ell q}|^{2}$, e.g., as $\sum_{q}(q^{2}{\mathsf{Var}}\,\widetilde{\sigma}_{\ell q})/\sum_{q}({\mathsf{Var}}\,\widetilde{\sigma}_{\ell q})=\Lambda_{\ell}^{-2}$ (substitute Eq. (49) into Eq. (53) and make use of the assumption that $\sigma_{\ell}(x)$ are mean-square differentiable stationary random processes, see G, so that the random spectral coefficients $\widetilde{\sigma}_{\ell q}$ are uncorrelated for different $q$). Since $\Lambda_{\ell}^{-1}\leq\Lambda^{-1}$ by the definition of $\Lambda$ and $\Lambda^{-1}\ll L_{\cal H}^{-1}$ by the above assumption, we obtain $\Lambda_{\ell}^{-1}\ll L_{\cal H}^{-1}$, which means that for any $\ell$, the function $H(\ell+q)$ changes little within the width of $\widetilde{\sigma}_{\ell q}$. As a result, the relation $H(\ell+q)\approx H(\ell)$ can be substituted into Eq. (60), yielding an approximation to the filter’s output $\varphi(x)$ under local stationarity:
$\varphi(x)\approx\utilde{\varphi}(x)=\sum\sigma_{\ell}(x)\,\widetilde{\alpha}_{l}\,H(l)\,\mathrm{e}^{\mathsf{i}\ell x}.$ (61)
(Its spherical counterpart is given in Eq. (8).) The approximation error in this equation is
$\delta\varphi(x)=\utilde{\varphi}(x)-\varphi(x)=\sum_{\ell}\widetilde{\alpha}_{\ell}\,\mathrm{e}^{\mathsf{i}\ell x}\sum_{q}\left[H(\ell+q)-H(\ell)\right]\,\widetilde{\sigma}_{\ell q}\,\mathrm{e}^{\mathsf{i}qx}.$ (62)
With $H(\ell+q)-H(\ell)\approx H^{\prime}(\ell)q=\kappa^{\prime}(\cdot)q/d\equiv\kappa^{\prime}(\cdot)qL_{\cal H}$, we have
$\operatorname{\mathbb{E}}[(\delta\varphi(x))^{2}\,|\,\sigma]\approx L_{\cal H}^{2}\sum_{\ell}[\kappa^{\prime}(\cdot)]^{2}\left|\sum_{q}q\,\widetilde{\sigma}_{\ell q}\,\mathrm{e}^{\mathsf{i}qx}\right|^{2}.$ (63)
Taking the expectation over the distribution of the spectral functions $\sigma_{\ell}(x)$ yields the mean error-variance
$\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}[(\delta\varphi(x))^{2}\,|\,\sigma]\approx L_{\cal H}^{2}\sum_{\ell}[\kappa^{\prime}(\cdot)]^{2}\sum_{q}q^{2}\operatorname{\mathbb{E}}|\widetilde{\sigma}_{\ell q}|^{2}=L_{\cal H}^{2}\sum_{\ell}[\kappa^{\prime}(\cdot)]^{2}\operatorname{\mathbb{E}}\left(\frac{\partial\sigma_{\ell}}{\partial x}\right)^{2},$ (64)
Denoting $\|\chi(x)\|^{2}=\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}(\chi^{2}(x){\mbox{$\,|\,$}}\sigma)$ (where $\chi(x)$ is a linearly transformed solution to our spatial model) and using Eqs. (53) and (55), we obtain the relative error-variance
$\frac{\|\delta\varphi\|^{2}}{\|\utilde{\varphi}\|^{2}}\leq\frac{\sum[\kappa^{\prime}(\cdot)]^{2}{\mathsf{Var}}\,\sigma_{\ell}}{\sum[\kappa(\cdot)]^{2}(\operatorname{\mathbb{E}}^{2}\sigma_{\ell}+{\mathsf{Var}}\,\sigma_{\ell})}\left(\frac{L_{\cal H}}{\Lambda}\right)^{2}\leq S\left(\frac{L_{\cal H}}{\Lambda}\right)^{2}\frac{\sum[\kappa^{\prime}(\cdot)]^{2}\operatorname{\mathbb{E}}^{2}\sigma_{\ell}}{\sum[\kappa(\cdot)]^{2}\operatorname{\mathbb{E}}^{2}\sigma_{\ell}}.$ (65)
Here, the rightmost fraction can be neglected since, by definition, $|\kappa(\cdot)|\leq 1$, $|\kappa^{\prime}(\cdot)|\leq 1$, and these functions have, roughly, the same widths so that $(\sum[\kappa^{\prime}(\cdot)]^{2}\operatorname{\mathbb{E}}^{2}\sigma_{\ell})/(\sum[\kappa(\cdot)]^{2}\operatorname{\mathbb{E}}^{2}\sigma_{\ell})\approx 1$. Therefore, the upper bound of the relative error standard deviation is, approximately,
$\frac{\|\delta\varphi\|}{\|\utilde{\varphi}\|}\leq\sqrt{S}\left(\frac{L_{\cal H}}{\Lambda}\right)\leq\left(\frac{L_{\cal H}}{\Lambda}\right).$ (66)
This means that the approximation error in Eq. (8) is small whenever the non-stationarity length scale is large enough, $L_{\cal H}\ll\Lambda$, or the process non-stationarity strength is low, $S\ll 1$ (with $\Lambda\geq L_{\cal H}$).
Applying the above technique to the variances $v={\mathsf{Var}}\,\varphi$ and $\utilde{v}={\mathsf{Var}}\,\utilde{\varphi}$ and exploiting the fact that for any $\ell$, $\operatorname{\mathbb{E}}|\widetilde{\sigma}_{\ell q}|^{2}$ is an even function of $q$, we obtain that their difference $\delta v=\utilde{v}-v$ satisfies

### v5

@@ Appendix D The most localized kernel in the stationary case
where it is straightforward to see that $\operatorname{\mathbb{E}}\widetilde{\xi}_{\ell m}\widetilde{\xi}_{\ell^{\prime}m^{\prime}}^{*}=\sigma_{\ell}^{2}\,\delta_{\ell}^{\ell^{\prime}}\delta_{m}^{m^{\prime}}$. The (real-valued) spectrum of the process is $f_{\ell}=\sigma_{\ell}^{2}$. The process $\xi_{\rm statio}(x)$ can be modeled as the convolution of the kernel
$u(\rho)=\sum^{\ell_{\rm max}}_{\ell=0}\frac{2\ell+1}{4\pi}\sigma_{\ell}P_{\ell}(\cos\rho)$ (53)
@@
Appendix E Identifiability of the model
@@
where $Q$ is the half-bandwidth of the processes $\sigma_{\ell}(x)$ as functions of location $x$. Note that since $\sigma_{-\ell}(x)=\sigma_{\ell}(x)$ (see C), we have $\widetilde{\sigma}_{-\ell,q}=\widetilde{\sigma}_{\ell q}$. Since $\sigma_{\ell}(x)$ is real-valued, we have $\widetilde{\sigma}_{\ell,-q}=\widetilde{\sigma}_{\ell q}^{*}$.
@@
Let us substitute the expansion Eq. (57) into Eq. (50) and change the summation variable $q$ to $n=\ell+q$666 In this Appendix (in contrast to the rest of the paper), we denote $\ell_{\rm max}$ by $L$ to simplify the notation. :
@@
so that we can uniquely restore $\widetilde{\sigma}_{Lq}$ for all $q$. To realize this, take $q=-Q$ in Eq. (60) (the lower right corner of the summation area in Fig. 9). We obtain $\widetilde{B}_{N,L-Q}=\widetilde{\sigma}_{LQ}\,\widetilde{\sigma}_{L,-Q}^{*}$. Recalling that $\widetilde{\sigma}_{\ell,-q}=\widetilde{\sigma}_{\ell q}^{*}$, we have $\widetilde{B}_{N,L-Q}=\widetilde{\sigma}_{LQ}^{2}$, which implies that $\widetilde{\sigma}_{LQ}$ is uniquely determined by $\widetilde{B}_{N,L-Q}$, up to sign. By the tight bandwidth assumption, $\widetilde{\sigma}_{LQ}\neq 0$, therefore, we recover $\widetilde{\sigma}_{Lq}$ for all $|q|<Q$ from Eq. (60) (the right edge of the summation parallelogram in Fig. 9), again, up to sign. The ambiguity in the sign disappears when it comes to $\widetilde{\sigma}_{L0}$, which is positive because $\sigma_{\ell}(x)>0$ due to the positive-definiteness constraint (see item (2) in section 3.8 and C). Thus, $\widetilde{\sigma}_{Lq}$ are uniquely recovered for all $q$ from the model covariances.
@@
Then, we consider $\widetilde{\xi}_{N-1}$ and realize that it, again, has just one term, $\widetilde{\alpha}_{L-1}\,\widetilde{\sigma}_{L-1,L-1+Q}$, besides the term that contains $\widetilde{\sigma}_{L,L+Q-1}$, which has already been recovered. This allows us to repeat the above process and recover $\widetilde{\sigma}_{L-1,L-1+q}$ for all $q\in[-Q,Q]$. And so on, we uniquely recover all non-zero $\widetilde{\sigma}_{\ell q}$, and therefore all spectral functions $\sigma_{\ell}(x)$, from the set of spectral covariances $\widetilde{B}_{n\bar{n}}$. This completes the proof that the non-stationary spatial convolution model on the circle is unique provided that the convolution kernel $u(x,\rho)$ is real-valued, is an even function of $\rho$, and the half-bandwidth $Q$ of $\sigma_{\ell}(x)$ is such that $\widetilde{\sigma}_{\ell Q}\neq 0$ for any $\ell$. On ${\mathbb{S}}^{2}$, the same reasoning is valid (not shown).
Appendix F Bandpass filtering of a locally stationary signal
Here we consider spatial bandpass filtering of a random process governed by the Locally Stationary Convolution Model. We consider the filtered process in the locally stationary regime with large $\Lambda$ or, formally, in the limit $\Lambda\to\infty$, with the mean process spectrum $\operatorname{\mathbb{E}}\sigma_{\ell}^{2}$ (see section 3.6) being kept constant. We examine the approximation errors involved in Eqs. (15) and Eq. (17) and show that these errors are small whenever $\Lambda$ is much greater than the the length scale of the filter’s impulse response function (formally, they vanish in the above limit). For simplicity of presentation, we consider the circular case.
@@ Appendix F Bandpass filtering of a locally stationary signal
$H(\ell)=\kappa\left(\frac{\ell-\ell^{\rm c}}{T}\right),$ (61)
where $\kappa(\cdot)$ is a smooth real-valued bell-shaped function of continuous argument such that $\kappa(0)=\max|\kappa(\cdot)|=1$ and $\max|\kappa^{\prime}(\cdot)|=1$, $\ell^{\rm c}$ is the band’s central wavenumber, and $T$ is the half-bandwidth (cf. Eq. (33)). Let $L_{\cal H}=1/T$ denote the length scale of the filter’s impulse response function. Apply the filter ${\cal H}$ to the process
@@
$\varphi(x)=\sum_{n=-\ell_{\rm max}-Q}^{\ell_{\rm max}+Q}H(n)\,\widetilde{\xi}_{n}\,\mathrm{e}^{\mathsf{i}nx}=\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\widetilde{\alpha}_{\ell}\,\mathrm{e}^{\mathsf{i}\ell x}\sum_{q=-Q}^{Q}H(\ell+q)\,\widetilde{\sigma}_{\ell q}\,\mathrm{e}^{\mathsf{i}qx},$ (63)
where, we recall, $\widetilde{\sigma}_{\ell q}$ are the spectral coefficients of the expansion $\sigma_{\ell}(x)=\sum_{q}\,\widetilde{\sigma}_{\ell q}\,\mathrm{e}^{\mathsf{i}qx}$. We suppose that the spatial filter in question is selected such that $L_{\cal H}\ll\Lambda$. Here $\Lambda$ is the non-stationarity length scale defined on the circle from